OpenStudy (anonymous):

Given the circle x^2+(y+1)^2-1=8, find its circle, radius and intercepts.

OpenStudy (anonymous):

add 1 to get\[x^2+(y+1)^2=9\]then the center is (0, -1) and the radius is 3. the y-intercepts are (0, -4) and (0, 2)

OpenStudy (anonymous):

and the x-intercepts are \[(2\sqrt{2}, 0)\] and \[(-2\sqrt{2}, 0)\]

OpenStudy (anonymous):

thanks for finishing the x-ints for me alsamixer!